what is the 10th,20th,and 30th terms of
1.)3,14,25,36,47
2.)25,38,51,64,77

Answer by Edwin McCravy(20056)

what is the 10th,20th,and 30th terms of
1.)3,14,25,36,47
2.)25,38,51,64,77
I'll just find the 10th term of 1.)  The others are done the same way.

Use the formula

an = a1 + (n-1)·d

You have a1 = 3, and n = 10. But you also need d.

To find d you subtract the 1st term from the 2nd term:

d = a2 - a1 = 14 - 3 = 11

But you should also check to make sure that the sequence is arithmetic,
otherwise the formula would not work:

d = a3 - a2 = 25 - 14 = 11
d = a4 - a3 = 36 - 25 = 11
d = a5 - a4 = 47 - 36 = 11

If any of those had been different, the above formula would not work.
But since they're all the same it will work.

For the 10th term, 

an = a1 + (n-1)·d
a10 = 3 + (10-1)·11
an = 3 + (9)·11
an = 3 + 99
an = 102

You can find all the others exactly the same way.

Edwin